OpenStudy (anonymous):

medal and fan to who ever helps me a = ?, b = 13, c = 21 A. 15.83 B. 15.87 C. 16.19 D. 16.49

OpenStudy (anonymous):

Is this a right triangle?

OpenStudy (anonymous):

there isnt a triangle

OpenStudy (anonymous):

Pythagorean Theorem. \(\sf a^2 + b^2 = c^2\) Plug in what we know: \(\sf a^2 + 13^2 = 21^2\) Solve for 'a'

OpenStudy (anonymous):